Consider that one mole of hydrogen gas is heated at a constant pressure from 340 K to 480 K. Calculate (a) the energy transferred by heat to the gas, (b) the increase in its internal energy and (c) the work done by the gas.

a

(a) 4.03 kJ; (b) 2.86 kJ; (c) 6.89 kJ

b

(a) 4.03 kJ; (b) 2.86 kJ; (c) 1.18 kJ

c

(a) 4.03 kJ; (b) 5.21 kJ; (c) 1.18 kJ

d

(a) 5.03 kJ; (b) 2.86 kJ; (c) 1.18 kJ
Answer #1

Given, 1 mole of hydrogen (H2) 2x (1.000) g = 2.016 OT 2 2 480-340 140 K and, specific heat of hydrogen gas (p = 14.3 5/g.k ((b) co/cv= 1.405 cv = 14.3/1.405 (cv= 10.18 Ilgik su = nCVOT = 2.016 x 10,18 x 140 AU 2.86 KJ (c) W: Q-OU 1.18 KJOption (b) is correct